# An Effective Deligne’s equidistribution theorem

### v1

@@ 4. Joint distribution
$n=n^{\prime}=1.$
The number of positive roots is also 1. So we have
$62(|R^{+}|+|R^{\prime+}|+n+n^{\prime}-1)=$ 6 .
A maximal compact subgroup of $\mathrm{SL}_{2}(\mathbf{C})$ is $\mathrm{SU}(2)$, and any element in $\mathrm{SU}(2)$ is conjugate to $(\begin{smallmatrix}e^{i\theta}&\\ &e^{-i\theta}\end{smallmatrix})$ for a unique $\theta\in[0,\pi]$. So $\mathrm{SU}(2)^{\natural}$ is identified with $[0,\pi]$. For any $x\in\mathbb{G}_{m}(\mathbf{F}_{q^{m}})$, the conjugacy class $\rho_{\mathrm{Kl}(1/2)}(F_{x})^{\mathrm{ss}}\in\mathrm{SU}(2)^{\natural}$ corresponds to the angle $\theta(x)\in[0,\pi]$, and $\rho_{\mathrm{Ai}(\alpha)}(F_{x})^{\mathrm{ss}}\in\mathrm{SU}(2)^{\natural}$ corresponds to the angle $\theta^{\prime}(x)\in[0,\pi]$. Using Weyl’s integration formula, one can verify $\mu_{\mathrm{SU}(2)^{\natural}}$ is identified with the Sato-Tate measure on $[0,\pi]$. Confer Proposition 2.1 below. Irreducible representations of $\mathrm{SL}_{2}$ are symmetric products of the standard representation. We claim that for any positive integers $k_{1},k_{2}$, $\mathrm{Sym}^{k_{1}}(\mathrm{Kl})\otimes\mathrm{Sym}^{k_{2}}(\mathrm{Ai})$ has no nonzero $\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})$-invariant. Our assertion then follows immediately from Theorem 1.7.

### v2

@@ 1. Introduction
$n=n^{\prime}=1,$
and the number of positive roots is
$|R^{+}|=|R^{\prime+}|=1.$
A maximal compact subgroup of $\mathrm{SL}_{2}(\mathbf{C})$ is $\mathrm{SU}(2)$, and any element in $\mathrm{SU}(2)$ is conjugate to $(\begin{smallmatrix}e^{i\theta}&\\ &e^{-i\theta}\end{smallmatrix})$ for a unique $\theta\in[0,\pi]$. So $\mathrm{SU}(2)^{\natural}$ is identified with $[0,\pi]$. For any $x\in\mathbb{G}_{m}(\mathbf{F}_{q^{m}})$, the conjugacy class $\rho_{\mathrm{Kl}(1/2)}(F_{x})^{\mathrm{ss}}\in\mathrm{SU}(2)^{\natural}$ corresponds to the angle $\theta(x)\in[0,\pi]$, and $\rho_{\mathrm{Ai}(\alpha)}(F_{x})^{\mathrm{ss}}\in\mathrm{SU}(2)^{\natural}$ corresponds to the angle $\theta^{\prime}(x)\in[0,\pi]$. Using Weyl’s integration formula, one can verify $\mu_{\mathrm{SU}(2)^{\natural}}$ is identified with the Sato-Tate measure on $[0,\pi]$. Confer Proposition 3.1 below. Irreducible representations of $\mathrm{SL}_{2}$ are symmetric products of the standard representation. We claim that for any positive integers $k_{1},k_{2}$, $\mathrm{Sym}^{k_{1}}(\mathrm{Kl})\otimes\mathrm{Sym}^{k_{2}}(\mathrm{Ai})$ has no nonzero $\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})$-invariant. Our assertion then follows from Theorem 1.7.